OpenStudy (anonymous):

Find standard form of the ellipse with a center of: (2,-1), and a vertex of:(2,8) I think that the minor axis of length is 5, but I'm not positive. I know that the formula is (X-h)^2/a^2 + (y-k)^2/b^2 = 1

OpenStudy (mertsj):

a = 9, a^2=81 If b = 5, b^2 = 25 This ellipse has major axis vertical so a^2 is under y^2

OpenStudy (mertsj):

\[\frac{(x-2)^2}{25}+\frac{(y+1)^2}{81}=1\]

OpenStudy (mertsj):

I don't know how you got that b = 5 but if it does, that is the equation.

OpenStudy (anonymous):

What would you do if it didn't equal five?

OpenStudy (mertsj):

You would have to have more information.

OpenStudy (anonymous):

Alright then, thank you so much for your help!

OpenStudy (mertsj):

Actually, if the length of the minor axis is 5 then b = 5/2 and b^2=25/4

OpenStudy (mertsj):

yw

OpenStudy (mertsj):

Does the problem give any additional information? Perhaps something about the foci or something?

OpenStudy (anonymous):

No. But I think that the minor axis of length is 5.

OpenStudy (mertsj):

Why do you think that?

OpenStudy (mertsj):

If you have posted the entire problem, there is nothing to suggest that.

OpenStudy (anonymous):

Because I copied down the problem from memory and that was part of the additional information.

OpenStudy (anonymous):

But I'm not positive.

OpenStudy (mertsj):

ok. Then 5 = 5/2

OpenStudy (mertsj):

Or maybe it gave the distance from the center to the co-vertex.

OpenStudy (mertsj):

Anyway, if you understand how to do the problem, the number isn't so important.

OpenStudy (anonymous):

Alright, I think I'm getting it. Thanks for your help!